- Press Release
- Oct 5, 2022
NOAA GOES SATOPS Morning Report 18 October 2000
SATOPS MORNING REPORT – October 18, 2000
No change in the status of the POES spacecraft.
NOAA-16 operations nominal.
POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
On F13/OPS47 Rev 28729, a 102.325 playback (4093 S/C) second Cat 1 data
loss was declared. SOCC crews could not establish a commanding link
with POGO ARTS using either the SOCC or PEAK ground equipment. The
crew tried to utilize Fairbanks CDA but, the site could not turn around
in time from a NOAA-14 support. Cause of linking problem is being
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
GOES-8 (East) Operations:
GOES-8 East/West maneuver #32 is scheduled for October 25 (Day 299)
at approximately 1658Z. The commanded thruster firing duration is
GOES-8 images cancelled for KOZ – 0409, 0439, and 0509Z.
GOES-10 (West) Operations:
GOES-10 East/West maneuver #14 is scheduled for October 24 (Day 298)
at approximately 1030Z. The commanded thruster firing duration is
GOES-10 images cancelled for KOZ: 0752, 0822 and 0852Z.
At 1323Z, 1415Z and 1740Z on Tuesday, October 17, GOES-10 frame breaks
were observed and attributed to high BER (unknown cause, but related to
the 13C antenna).
At 1513Z on Tuesday, October 17, a GOES-10 frame break were observed
and attributed to loss of signal synch.
At 1740Z on Tuesday, October 17, GOES-10 support was transferred to
the 16M HR2 antenna. No frame breaks have been recorded subsequent
to the transfer.
Storage Mode Spacecraft Operations:
GOES-9 and GOES-11 remain in the Z-Axis Precession (ZAP) storage mode
at approximately 105 deg W.
On Tuesday, October 17, GOES-9 support was successfully shifted from
the HR2 antenna to the 13M ‘C’ antenna (via the 14.2M).
GOES operations were nominal over the past 24 hours.
Questions or comments to firstname.lastname@example.org